# PDV: Prompt Directional Vectors for Zero-shot Composed Image Retrieval

### v1

@@ 4.3 Effect of Using the PDV
In this section we explore the impact of the three proposed uses of the prompt direction vector (PDV): Using the PDV to augment text queries (PDV-T, see Sec. 4.3.1), using the PDV to augment image queries (PDV-I, see Sec. 4.3.2), and using the PDV in queries that fuse image and text data (PDV-F, see Sec. 4.3.3).
4.3.1 Analysing the PDV for Text (PDV-T)
To investigate how scaling the prompt vector, $\Delta_{PDV}$, affects retrieval performance with composed text embeddings, we conducted experiments using two zero-shot approaches (CIReVL and Pic2Word) with different backbone networks across three datasets. We evaluated the performance by varying the scaling parameter, $\alpha$ (Eq. 3), from -0.5 to 3 by an interval of 0.1.
The results are presented in Figure 3(a). To account for scale variations across different experiments, we report relative recall values, where a baseline of zero is established at $\alpha=1$. As shown in Figure 3(a), varying $\alpha$ leads to significant changes in relative recall performance [^3] See supplementary material for Recall@10 and Recall@50 figures. Our analysis reveals method-specific patterns across datasets. With CIReVL, increasing $\alpha$ improves relative recall on both FashionIQ and CIRCO datasets. In contrast, Pic2Word shows no significant improvement on FashionIQ and CIRR when varying $\alpha$, while CIRCO’s performance improves when $\alpha$ is reduced to 0.8-1.0. This divergent behavior is fundamentally linked to each method’s ability to generate an accurate $\Delta_{PDV}$. As demonstrated in Tables 2 and 1, CIReVL consistently outperforms Pic2Word across various benchmarks, indicating its superior ability to generate a more accuraute composed query, and thus a more accurate $\Delta_{PDV}$. Consequently, increasing $\alpha$ yields greater benefits for CIReVL compared to Pic2Word.
4.3.2 Analysing the PDV for Image (PDV-I)
To evaluate whether $\Delta_{PDV}$ enhances the retrieval performance of image embeddings, we conducted experiments following the protocol described in Section 4.3.1. We modified image embeddings by adding $\Delta_{PDV}$ scaled with $\alpha$ values ranging from -0.5 to 2.0, where $\alpha=0$ represents the original image-only embeddings. As shown in Figure 3(b), Recall@K exhibits a positive correlation with $\alpha$ for values below 1. This upward trend continues until $\alpha=2.0$ for CIReVL, while Pic2Word’s performance peaks when $\alpha$ reaches 1.4.
Following the methodology in Section 4.3.1, we conduct similar visualizations, with results shown in Figure 4b. As with PDV-T, increasing $\alpha$ leads to stronger alignment between retrieved results and the prompt. When $\alpha$ exceeds 0.5, the results exhibit semantic relationships to the query, while $\alpha$ values below 0.5 yield results opposing the prompt’s intent. Notably, PDV-I’s top retrievals demonstrate higher visual similarity to reference images compared to PDV-F, as evidenced by the preserved design elements in the clothing item (left) and laptop (middle). This characteristic is particularly valuable for applications include fashion search [21] and logo retrieval [19], where visual similarity plays a crucial role.
Figure 4: Visualisation of the impact of $\alpha$/$\beta$ scaling to top-5 retrieval results. CIReVL with ViT-B-32 Clip model is the baseline method used. Representative examples with prompts from three datasets: FashionIQ (left), CIRR (middle), and CIRCO (right) are shown at the top. Green and blue bounding boxes indicate true positives and near-true positives, respectively.
4.3.3 Analysing PDV Fusion (PDV-F)
Finally, we evaluate the effectiveness of fusing image and text-composed embeddings by varying the fusion parameter, $\beta$, from 0 to 1. At $\beta=0$, the model relies solely on composed image embeddings, while at $\beta=1$, it uses only composed text embeddings. As shown in Figure 3(c), the fusion of both embeddings consistently outperforms either embedding type alone. Optimal retrieval performance is typically achieved when $\beta$ is between 0.4 and 0.8.
@@ 4.3.3 Analysing PDV Fusion (PDV-F)
Arch Method k=5 k=10 k=25 k=50 k=1 k=5 k=10 k=50 k=1 k=2 k=3
ViT-B/32 Image-only † 1.34 1.60 2.12 2.41 6.89 22.99 33.68 59.23 21.04 41.04 60.31
Text-only † 2.56 2.67 2.98 3.18 21.81 45.22 57.42 81.01 62.24 81.13 90.70
Image + Text † 2.65 3.25 4.14 4.54 11.71 35.06 48.94 77.49 32.77 56.89 74.96
PALAVRA † 4.61 5.32 6.33 6.80 16.62 43.49 58.51 83.95 41.61 65.30 80.94
SEARLE † 9.35 9.94 11.13 11.84 24.00 53.42 66.82 89.78 54.89 76.60 88.19
CIReVL † 14.94 15.42 17.00 17.82 23.94 52.51 66.0 86.95 60.17 80.05 90.19
CIReVL + PDV-F 15.26 16.19 18.16 19.10 31.93 64.19 75.16 92.58 61.86 81.11 90.84
ViT-L/14 Pic2Word 6.81 7.49 8.51 9.07 23.69 51.32 63.66 86.21 53.61 74.34 87.28
Pic2Word + PDV-F 7.77 8.70 9.80 10.42 23.81 51.86 64.34 86.84 52.58 73.54 86.89
SEARLE † 11.68 12.73 14.33 15.12 24.24 52.48 66.29 88.84 53.76 75.01 88.19
CIReVL † 18.57 19.01 20.89 21.80 24.55 52.31 64.92 86.34 59.54 79.88 89.69
CIReVL + PDV-F 21.19 22.49 24.67 25.74 34.92 66.58 76.27 92.94 63.95 83.11 92.02
ViT-G/14 CIReVL † 26.77 27.59 29.96 31.03 34.65 64.29 75.06 91.66 67.95 84.87 93.21
CIReVL + PDV-F 25.24 26.80 29.47 30.56 36.75 67.54 76.99 93.18 65.06 82.53 91.74
@@
Table 1: Performance comparison on CIRCO and CIRR test datasets across different backbone architectures. For CIRCO, mAP@k is reported, while for CIRR both Recall@k and Rs@k metrics are used. †denotes that numbers are taken from the original paper.
@@ 4.4 ZS-CIR Benchmark Comparison
Backbone Method R@10 R@50
ViT-B/32 Image-only † 5.90 13.37
Text-only † 18.70 36.84
Image + Text 14.78 29.60
PALAVRA † 19.76 37.25
SEARLE † 22.89 42.53
CIReVL † 28.29 49.35
CIReVL + PDV-F 33.82 54.94
ViT-L/14 Pic2Word † 24.70 43.70
Pic2Word + PDV-F 26.12 45.26
SEARLE † 25.56 46.23
CIReVL † 28.55 48.57
CIReVL + PDV-F 36.78 56.82
ViT-G/14 CIReVL † 32.19 52.36
CIReVL + PDV-F 41.72 61.39
Table 2: Comparison of average recall values of different methods on Fashion-IQ validation dataset using various backbone models. †denotes that numbers are taken from the original paper.
We evaluated PDV-F in combination with two baseline approaches (CIReVL and Pic2Word) across three benchmarks, setting $\beta$ to 0.7 for CIReVL and 0.9 for Pic2Word based on the fusion ablation results shown in Figure 3(c). The results are presented in Tables 1 and 2. On the FashionIQ benchmark, PDV-F yields substantial improvements for both baseline approaches, with CIReVL showing particularly strong gains that scale with backbone size. Similarly, both methods demonstrate significant performance improvements on CIRCO and CIRR datasets. Notably, CIReVL achieves larger improvements compared to Pic2Word, with the most substantial gains observed when using small and medium backbone architectures.
@@ Appendix A Appendix
In this section, we present additional quantitative results.
A.0.1 Ablation Analysis
While Figure 3 in the main paper illustrates the effects of scaling factor $\alpha$ and fusion factor $\beta$ on Recall@5 performance across various PDV applications, Figures 5, 6, and 7 present complementary results for Recall@10 and Recall@50 metrics.
@@ A.0.1 Ablation Analysis
A.0.2 ZS-CIR Benchmark Comparison
We also provide additional results achieved on the test set of CIRCO and CIRR datasets (see Table 3), and the validation set of FashionIQ (see Table 4).
  • We evaluate PDV integrated with SEARLE [2], a variant of Pic2Word, but more focussed on the salient object in the given image. In this implementation, PDV-F’s $\beta$ parameter is set to 0.8 across all three datasets, though ablation studies (Figure 7) suggest this value may not be optimal for all scenarios.
  • We provide detailed performance breakdowns for individual sub-datasets within FashionIQ in Table 4.
  • The optimal performance achieved by both PDV-I and PDV-F on the FashionIQ validation datasets is presented in Table 4.
PDV-T and PDV-F enhance the performance of all three baseline approaches on the FashionIQ dataset. While PDV-I does not improve upon text-inversion methods like Pic2Word and SEARLE, its results significantly outperform image-only baselines and contribute to fusion results that exceed baseline performance.
In contrast, PDV-F demonstrates improved performance over CIReVL and Pic2Word on both CIRCO and CIRR datasets, though it fails to enhance SEARLE’s results. We hypothesize this limitation stems from SEARLE’s single object oriented text inversion approach, which may be less effective for the CIRCO and CIRR datasets which contain multiple objects and complex scenes. This characteristic explains why PDV-F successfully improves the performance of SEARLE in FashionIQ, which features simpler compositions, but not in the CIRCO and CIRR datasets.
@@ A.0.2 ZS-CIR Benchmark Comparison
Dataset CIRCO CIRR
Metric mAP@k Recall@k $R_{s}$@k
Arch Method k=5 k=10 k=25 k=50 k=1 k=5 k=10 k=50 k=1 k=2 k=3
ViT-B/32 Image-only † 1.34 1.60 2.12 2.41 6.89 22.99 33.68 59.23 21.04 41.04 60.31
Text-only † 2.56 2.67 2.98 3.18 21.81 45.22 57.42 81.01 62.24 81.13 90.70
Image + Text † 2.65 3.25 4.14 4.54 11.71 35.06 48.94 77.49 32.77 56.89 74.96
PALAVRA † 4.61 5.32 6.33 6.80 16.62 43.49 58.51 83.95 41.61 65.30 80.94
SEARLE † 9.35 9.94 11.13 11.84 24.00 53.42 66.82 89.78 54.89 76.60 88.19
SEARLE + PDV-F 7.26 8.12 9.4 10.12 22.0 51.40 65.18 88.92 50.17 72.75 88.92
CIReVL † 14.94 15.42 17.00 17.82 23.94 52.51 66.0 86.95 60.17 80.05 90.19
CIReVL + PDV-F 15.26 16.19 18.16 19.10 31.93 64.19 75.16 92.58 61.86 81.11 90.84
ViT-L/14 Pic2Word 6.81 7.49 8.51 9.07 23.69 51.32 63.66 86.21 53.61 74.34 87.28
Pic2Word + PDV-F 7.77 8.70 9.80 10.42 23.81 51.86 64.34 86.84 52.58 73.54 86.89
SEARLE † 11.68 12.73 14.33 15.12 24.24 52.48 66.29 88.84 53.76 75.01 88.19
SEARLE + PDV-F 10.25 11.53 13.5 14.32 23.08 51.33 64.70 88.34 50.27 72.70 86.05
CIReVL † 18.57 19.01 20.89 21.80 24.55 52.31 64.92 86.34 59.54 79.88 89.69
CIReVL + PDV-F 21.19 22.49 24.67 25.74 34.92 66.58 76.27 92.94 63.95 83.11 92.02
ViT-G/14 CIReVL † 26.77 27.59 29.96 31.03 34.65 64.29 75.06 91.66 67.95 84.87 93.21
CIReVL + PDV-F 25.24 26.80 29.47 30.56 36.75 67.54 76.99 93.18 65.06 82.53 91.74
Table 3: Performance comparison on the CIRCO and CIRR test datasets across different backbone architectures. For CIRCO, mAP@k is reported, while for CIRR both Recall@k and Rs@k metrics are used. †denotes that numbers are taken from the original paper.
Fashion-IQ Shirt Dress Toptee Average
Backbone Method R@10 R@50 R@10 R@50 R@10 R@50 R@10 R@50
ViT-B/32 Image-only † 6.92 14.23 4.46 12.19 6.32 13.77 5.90 13.37
Text-only † 19.87 34.99 15.42 35.05 20.81 40.49 18.70 36.84
Image + Text † 13.44 26.25 13.83 30.88 17.08 31.67 14.78 29.60
PALAVRA † 21.49 37.05 17.25 35.94 20.55 38.76 19.76 37.25
SEARLE 24.14 41.81 18.39 38.08 25.91 47.02 22.81 42.30
SEARLE + PDV-T 24.48 42.30 18.79 38.47 25.91 47.02 23.03 42.60
SEARLE + PDV-I 18.25 32.14 18.74 39.17 21.77 38.45 19.59 36.59
SEARLE + PDV-F 24.53 41.17 21.52 41.89 25.80 46.46 23.95 43.17
CIReVL † 28.36 47.84 25.29 46.36 31.21 53.85 28.29 49.35
CIReVL + PV-T 29.34 48.92 26.23 48.24 34.57 57.52 30.04 51.56
CIReVL + PV-I 28.95 45.88 29.00 49.13 34.22 56.09 30.72 50.37
CIReVL + PV-F 32.04 51.03 32.82 54.88 36.61 58.90 33.82 54.94
ViT-L/14 Pic2Word 25.96 43.52 19.63 40.90 27.28 47.83 24.29 44.08
Pic2Word + PV-T 26.10 43.52 19.88 41.10 27.59 47.83 24.52 44.15
Pic2Word + PV-I 24.39 39.89 18.69 36.79 25.80 43.04 22.96 39.91
Pic2Word + PV-F 28.21 44.55 21.12 42.34 29.02 48.90 26.12 45.26
SEARLE 26.84 45.19 20.08 42.19 28.40 49.62 25.11 45.67
SEARLE + PDV-T 27.77 46.22 20.67 43.08 28.67 49.77 25.70 46.36
SEARLE + PDV-I 23.50 38.71 22.95 44.03 27.43 45.49 24.63 42.74
SEARLE +PDV-F 28.95 46.86 23.50 46.11 31.26 52.07 27.90 48.35
CIReVL † 29.49 47.40 24.79 44.76 31.36 53.65 28.55 48.57
CIReVL + PV-T 32.14 50.39 26.33 47.84 34.93 57.88 31.13 52.04
CIReVL + PV-I 33.95 50.39 31.73 53.00 39.16 59.95 34.95 54.45
CIReVL + PV-F 36.85 53.68 33.32 55.38 40.18 61.40 36.78 56.82
ViT-G/14* CIReVL † 33.71 51.42 27.07 49.53 35.80 56.14 32.19 52.36
CIReVL + PV-T 35.57 52.45 30.34 53.00 39.42 60.43 35.11 55.29
CIReVL + PV-I 37.78 54.32 35.85 58.25 43.86 63.95 39.16 58.84
CIReVL + PV-F 41.51 58.10 38.32 61.87 45.33 64.20 41.72 61.39
Table 4: Comparison of average recall values of different methods on Fashion-IQ validation dataset using various backbone models. †denotes that numbers are taken from the original paper..

### v2

@@ 4.1 Effect of Using the PDV
We now explore the impact of the three proposed uses of the PDV: Using the PDV to augment text queries (PDV-T, see Sec. 2), using the PDV to augment image queries (PDV-I, see Sec. 2), and using the PDV in queries that fuse image and text data (PDV-F, see Sec. 4).
Fashion-IQ Shirt Dress Toptee Average
Backbone Method $\beta$ $\alpha_{I}$ $\alpha_{T}$ R@10 R@50 R@10 R@50 R@10 R@50 R@10 R@50
ViT-B/32 SEARLE - - - 24.14 41.81 18.39 38.08 25.91 47.02 22.81 42.30
SEARLE + PDV-F 0.9 1.1 0.9 24.83 41.71 20.13 41.40 25.96 47.17 23.64 43.43
CIReVL † - - - 28.36 47.84 25.29 46.36 31.21 53.85 28.29 49.35
CIReVL + PDV-F 0.75 1.4 1.4 32.88 52.80 32.67 54.49 38.91 61.81 34.82 56.37
LDRE † - - - 27.38 46.27 19.97 41.84 27.07 48.78 24.81 45.63
SEIZE † - - - 29.38 47.97 25.37 46.84 32.07 54.78 28.94 49.86
ViT-L/14 Pic2Word 25.96 43.52 19.63 40.90 27.28 47.83 24.29 44.08
Pic2Word + PV-F 0.8 1.0 1.0 28.21 44.55 20.92 42.24 29.02 48.90 26.05 45.23
SEARLE - - - 26.84 45.19 20.08 42.19 28.40 49.62 25.11 45.67
SEARLE +PDV-F 0.8 1.2 1.0 28.66 46.76 23.60 46.41 31.00 52.32 27.75 48.50
CIReVL † 29.49 47.40 24.79 44.76 31.36 53.65 28.55 48.57
CIReVL + PDV-F 0.55 1 1.3 37.78 54.22 33.61 56.07 41.61 62.16 37.67 57.48
LinCIR - - - 29.10 46.81 20.92 42.44 28.81 50.18 26.82 46.49
SEIZE - - - 33.04 53.22 30.93 50.76 35.57 58.64 33.18 54.21
ViT-G/14 Pic2Word - - - 33.17 50.39 25.43 47.65 35.24 57.62 31.28 51.89
SEARLE - - - 36.46 55.35 28.16 50.32 39.83 61.45 34.81 55.71
CIReVL † - - - 33.71 51.42 27.07 49.53 35.80 56.14 32.19 52.36
CIReVL + PV-F 0.6 1.4 1.4 41.90 58.19 40.70 62.82 48.09 67.77 43.56 62.93
LinCIR - - - 46.76 65.11 38.08 60.88 50.48 71.09 45.11 65.69
SEIZE - - - 43.60 65.42 39.61 61.02 45.94 71.12 43.05 65.85
Table 1: Average recall for different methods on Fashion-IQ validation dataset. † denotes that numbers are taken from the original paper.
@@
Arch Method $\beta$ $\alpha_{I}$ $\alpha_{T}$ k=5 k=10 k=25 k=50 k=1 k=5 k=10 k=50 k=1 k=2 k=3
ViT-B/32 PALAVRA[6] - - - 4.61 5.32 6.33 6.80 16.62 43.49 58.51 83.95 41.61 65.30 80.94
SEARLE † - - - 9.35 9.94 11.13 11.84 24.00 53.42 66.82 89.78 54.89 76.60 88.19
SEARLE + PDV-F 0.9 1.4 1.2 9.99 10.50 11.70 12.40 24.53 53.71 67.33 89.81 56.94 78.05 88.99
CIReVL † - - - 14.94 15.42 17.00 17.82 23.94 52.51 66.00 86.95 60.17 80.05 90.19
CIReVL + PDV-F 0.75 1.4 1.2 19.90 20.61 22.64 23.52 33.25 64.15 75.23 92.43 65.81 83.76 92.10
LDRE - - - 17.81 18.04 19.73 20.67 25.69 55.52 68.77 89.86 60.10 80.58 91.04
LDRE + PDV-F 0.75 1.4 1.4 17.80 18.78 20.61 21.56 29.30 60.39 72.51 91.42 63.06 82.36 91.54
SEIZE - - - 19.04 19.64 21.55 22.49 27.47 57.42 70.17 - 65.59 84.48 92.77
ViT-L/14 Pic2Word - - - 6.81 7.49 8.51 9.07 23.69 51.32 63.66 86.21 53.61 74.34 87.28
Pic2Word + PDV-F 0.85 1.2 1.0 7.74 8.67 9.77 10.37 23.90 51.95 64.63 87.04 53.16 74.07 87.08
SEARLE † - - - 11.68 12.73 14.33 15.12 24.24 52.48 66.29 88.84 53.76 75.01 88.19
SEARLE + PDV-F 0.85 1.4 1.2 12.58 13.57 15.30 16.07 25.64 53.61 66.58 88.55 55.83 76.48 88.53
CIReVL † - - - 18.57 19.01 20.89 21.80 24.55 52.31 64.92 86.34 59.54 79.88 89.69
CIReVL + PDV-F 0.75 1.4 1.2 25.67 26.61 28.81 29.95 36.24 66.17 76.96 92.29 68.07 85.35 93.47
LDRE - - - 22.32 23.75 25.97 27.03 26.68 55.45 67.49 88.65 60.39 80.53 90.15
LDRE + PDV-F 0.75 1.4 1.4 25.23 26.52 28.94 29.95 30.16 59.98 71.90 90.87 63.66 82.87 91.57
LinCIR - - - 12.59 13.58 15.00 15.85 25.04 53.25 66.68 - 57.11 77.37 88.89
SEIZE - - - 24.98 25.82 28.24 29.35 28.65 57.16 69.23 - 66.22 84.05 92.34
ViT-G/14 CIReVL † - - - 26.77 27.59 29.96 31.03 34.65 64.29 75.06 91.66 67.95 84.87 93.21
CIReVL + PDV-F 0.75 1.4 1.2 30.02 31.46 34.01 35.08 38.15 67.93 77.90 92.77 69.37 85.37 93.45
LDRE - - - 33.30 34.32 37.17 38.27 37.40 66.96 78.17 93.66 68.84 85.64 93.90
LDRE + PDV-F 0.75 1.4 1.4 34.88 36.41 39.12 40.23 42.51 72.22 81.71 94.94 72.39 88.34 94.80
SEARLE - - - 13.20 13.85 15.32 16.04 34.80 64.07 75.11 - 68.72 84.70 93.23
LinCIR - - - 19.71 21.01 23.13 24.18 35.25 64.72 76.05 - 63.35 82.22 91.98
SEIZE - - - 32.46 33.77 36.46 37.55 38.87 69.42 79.42 - 74.15 89.23 95.71
@@
Table 2: Performance comparison on CIRCO and CIRR test datasets. As in previous works, for CIRCO, mAP@k is reported, while for CIRR both Recall@k and $R_{s}$@k metrics are used. † denotes that numbers are taken from the original paper.
Analysing the PDV for Text (PDV-T) To investigate how scaling the prompt vector, $\Delta_{PDV}$, affects retrieval performance with composed text embeddings, we conducted experiments using two zero-shot approaches (CIReVL and Pic2Word) with different backbone networks across three datasets. We evaluated the performance by varying the scaling parameter, $\alpha$ (Eq. 3), from -0.5 to 3 by an interval of 0.1.
The results are presented in Figure 3(a). To account for scale variations across different experiments, we report relative recall values, where a baseline of zero is established at $\alpha=1$. As shown in Figure 3(a), varying $\alpha$ leads to significant changes in relative recall performance [^5] See supplementary material for Recall@10 and Recall@50 figures. Our analysis reveals method-specific patterns across datasets. With CIReVL, increasing $\alpha$ improves relative recall on both FashionIQ and CIRCO datasets. In contrast, Pic2Word shows no significant improvement on FashionIQ and CIRR when varying $\alpha$, while CIRCO’s performance improves when $\alpha$ is reduced to 0.8-1.0. This divergent behavior is fundamentally linked to each method’s ability to generate an accurate $\Delta_{PDV}$. As demonstrated in Tables 1 and 2, CIReVL consistently outperforms Pic2Word across various benchmarks, indicating its superior ability to generate a more accuraute composed query, and thus a more accurate $\Delta_{PDV}$. Consequently, increasing $\alpha$ yields greater benefits for CIReVL compared to Pic2Word.
Analysing the PDV for Image (PDV-I) To evaluate whether $\Delta_{PDV}$ enhances the retrieval performance of image embeddings, we conducted experiments following the protocol described in Section 2. We modified image embeddings by adding $\Delta_{PDV}$ scaled with $\alpha$ values ranging from -0.5 to 2.0, where $\alpha=0$ represents the original image-only embeddings. As shown in Figure 3(b), Recall@K exhibits a positive correlation with $\alpha$ for values below 1. This upward trend continues until $\alpha=2.0$ for CIReVL, while Pic2Word’s performance peaks when $\alpha$ reaches 1.4. The performance of PDV-I was evaluated on the CIRR and CIRCO datasets by comparing it with other visual embedding-based methods, as detailed in Table 3. The results reveal that PDV-I achieved marginal improvements over existing approaches.
Following the methodology in Section 2, we conduct similar visualizations, with results shown in Figure 4b. As with PDV-T, increasing $\alpha$ leads to stronger alignment between retrieved results and the prompt. When $\alpha$ exceeds 0.5, the results exhibit semantic relationships to the query, while $\alpha$ values below 0.5 yield results opposing the prompt’s intent. Notably, PDV-I’s top retrievals demonstrate higher visual similarity to reference images compared to PDV-F, as evidenced by the preserved design elements in the clothing item (left) and laptop (middle). This characteristic is particularly valuable for applications include fashion search [22] and logo retrieval [20], where visual similarity plays a crucial role.
Figure 4: Visualisation of the impact of $\alpha$/$\beta$ scaling on top-5 retrieval results. CIReVL with ViT-B-32 Clip model is the baseline method used. Representative examples with prompts from three datasets: FashionIQ (left), CIRR (middle), and CIRCO (right) are shown at the top. Green and blue bounding boxes indicate true positives and near-true positives, respectively.
Analysing PDV Fusion (PDV-F) Finally, we evaluate the effectiveness of fusing image and text-composed embeddings by varying the fusion parameter, $\beta$, from 0 to 1 while maintaining $\alpha=1$ for both PDV-I and PDV-F. At $\beta=0$, the model relies solely on composed image embeddings, while at $\beta=1$, it uses only composed text embeddings. As shown in Figure 3(c), the fusion of both embeddings consistently outperforms using either embedding type alone. Optimal retrieval performance is typically achieved when $\beta$ is between 0.4 and 0.8.
@@ 4.2 ZS-CIR Benchmark Comparison
Dataset CIRCO CIRR
Metric mAP@k Recall@k $R_{s}$@k
Arch Method $\alpha_{I}$ k=5 k=10 k=25 k=50 k=1 k=5 k=10 k=50 k=1 k=2 k=3
ViT-B/32 Image-only † - 1.34 1.60 2.12 2.41 6.89 22.99 33.68 59.23 21.04 41.04 60.31
Text-only † - 2.56 2.67 2.98 3.18 21.81 45.22 57.42 81.01 62.24 81.13 90.70
Image + Text † - 2.65 3.25 4.14 4.54 11.71 35.06 48.94 77.49 32.77 56.89 74.96
SEARLE + PDV-I 1.5 4.77 5.23 6.31 6.82 16.65 42.53 55.16 81.42 44.68 67.78 82.94
CIReVL + PDV-I 2.0 10.29 10.80 12.23 12.93 27.18 56.53 67.76 87.64 59.81 79.59 90.15
LDRE + PDV-I 2.0 8.00 8.88 10.06 10.72 23.37 51.21 63.69 85.57 55.57 76.63 88.15
Table 3: PDV-I performance on CIRCO and CIRR test datasets. Note that the image-only approach utilizes the visual embedding of the reference image, whereas the text-only approach employs the text embedding of the prompt.
We evaluated PDV-F alongside four baseline approaches (CIReVL, LDRE, Pic2Word, and SEARLE) across three benchmarks. Notably, CIReVL was tested with three different backbones on three datasets, as its models and intermediate results are publicly available. However, for the remaining methods, we conducted partial evaluations due to limited open-source availability or restricted support.
The numerical results are presented in Tables 1 and 2. On the FashionIQ benchmark, PDV-F yields substantial improvements for all baseline approaches, with CIReVL showing particularly strong gains that scale with backbone size. Similarly, all methods demonstrate significant performance improvements on CIRCO and CIRR datasets. Notably, CIReVL achieves larger improvements compared to other methods, with the most substantial gains observed when using small and medium backbone architectures. Our PDV-F implementation within the CIReVL framework consistently outperformed other state-of-the-art methods, including LinCIR and SEIZE, across most evaluation metrics. Similar to SEIZE, PDV-F offers the advantage of being entirely training-free; however, unlike SEIZE, it does not significantly increase feature extraction computational costs. While LinCIR demonstrates exceptional inference speed, it lacks the training-free nature of our approach, requiring dedicated model training before deployment.
@@ Appendix A Appendix
A.1 PDV Algorithm and Code
The PDV algorithm is given in Algorithm 1, and the code is shown in Figure 5. The implementation of PDV is very intuitive, and it could be easily integrated with any ZS-CIR approaches.
Algorithm 1 Calculate PDV Features
1:function CalculatePDVFeatures($\mathbf{f}_{\text{text}}$, $\mathbf{f}_{\text{text\_composed}}$, $\mathbf{f}_{\text{image}}$, $\alpha_{i}$, $\alpha_{t}$, $\beta$)
2:     $\mathbf{f}_{\text{text}}\leftarrow\text{normalize}(\mathbf{f}_{\text{text}})$
3:     $\mathbf{f}_{\text{text\_composed}}\leftarrow\text{normalize}(\mathbf{f}_{\text{text\_composed}})$
4:     $\mathbf{f}_{\text{image}}\leftarrow\text{normalize}(\mathbf{f}_{\text{image}})$
5:     $\mathbf{pdv}\leftarrow\mathbf{f}_{\text{text\_composed}}-\mathbf{f}_{\text{text}}$
6:     $\mathbf{f}_{\text{PDVI}}\leftarrow\mathbf{f}_{\text{image}}+\alpha_{i}\cdot\mathbf{pdv}$
7:     $\mathbf{f}_{\text{PDVT}}\leftarrow\mathbf{f}_{\text{text}}+\alpha_{t}\cdot\mathbf{pdv}$
8:     $\mathbf{f}_{\text{PDVF}}\leftarrow(1-\beta)\cdot\mathbf{f}_{\text{PDVI}}+\beta\cdot\mathbf{f}_{\text{PDVT}}$
9:     return $\text{normalize}(\mathbf{f}_{\text{PDVF}})$
10:end function
Figure 5: Python function for calculating PDV features.
A.2 Additional Results
A.2.1 Additional Results
We also provide additional PDV-I results achieved on the validation set of the FashionIQ dataset, as shown in Table 4. PDV-I also achieved significant improvements over existing approaches that directly leverage image embeddings for retrieval.
Lastly, we provide a detailed visualization of the impact of $\alpha$/$\beta$ scaling on top-5 retrieval results. Figure 6 illustrates the performance of CIReVL with the ViT-B-32 CLIP model across three different datasets.
Fashion-IQ Shirt Dress Toptee Average
Backbone Method $\alpha$ R@10 R@50 R@10 R@50 R@10 R@50 R@10 R@50
ViT-B/32 Image-only † - 6.92 14.23 4.46 12.19 6.32 13.77 5.90 13.37
Text-only † - 19.87 34.99 15.42 35.05 20.81 40.49 18.70 36.84
Image + Text † - 13.44 26.25 13.83 30.88 17.08 31.67 14.78 29.60
SEARLE + PDV-I 24.48 42.30 18.79 38.47 25.91 47.02 23.03 42.60
CIReVL + PDV-I 28.95 45.88 29.00 49.13 34.22 56.09 30.72 50.37
Table 4: PDV-I performance on FashionIQ val datasets. † denotes that numbers are taken from the original paper.
A.2.2 Ablation Analysis
While Figure 3 in the main paper illustrates the effects of scaling factor $\alpha$ and fusion factor $\beta$ on Recall@5 performance across various PDV applications, Figures 7, 8, and 9 present complementary results for Recall@10 and Recall@50 metrics.
@@ A.2.2 Ablation Analysis
Figure 6: Visualisation of the impact of $\alpha$/$\beta$ scaling on top-5 retrieval results. CIReVL with ViT-B-32 Clip model is the baseline method used. Representative examples with prompts from three datasets: FashionIQ (left), CIRR (middle), and CIRCO (right) are shown at the top. Green and blue bounding boxes indicate true positives and near-true positives, respectively.